OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (anonymous):

that formula cos A = .....

OpenStudy (anonymous):

\[\cos \theta = \frac{ u + v }{ u \times v }\]

OpenStudy (anonymous):

yes yes :)

OpenStudy (anonymous):

Its odd though, when I try to apply it - it doesn't seem to come out properly. For instance, I do -5 multiplied by -4 which gives me 20. -4 times -3 which gives me 12, I add those together and get 32 on the numerator.

OpenStudy (anonymous):

Then on the denominator, I'm stuck because in an example it tells me to square root u and v.

OpenStudy (anonymous):

but its like u1v1 +u2v1 /

OpenStudy (anonymous):

Aaaaaah.

OpenStudy (anonymous):

I was doing it the other way around.

OpenStudy (anonymous):

wait i will use the equation?

OpenStudy (anonymous):

\[\frac{ u_1v_1+u_2v_2 }{ \sqrt{u_1^2u_2^2}.\sqrt{v_1^2v_2^2} }\]

OpenStudy (anonymous):

Okay, so you don't do u1 and u2, you do u1 and v1..

OpenStudy (anonymous):

yeah we do that

OpenStudy (anonymous):

we do them

OpenStudy (anonymous):

and we get cos = something something

OpenStudy (anonymous):

and from that we find the answer

OpenStudy (anonymous):

Yes, but that is where I'm stuck. I get the numbers at the bottom but when I try to figure it out my calculator says Domain Error.

OpenStudy (anonymous):

hah wait i will solve

OpenStudy (anonymous):

top is 32

OpenStudy (anonymous):

I see.. The answer choices however are: -9.1 degrees, 1.8 degrees, 0.9 degrees, and 11.8 degrees.

OpenStudy (anonymous):

nooo >_> i did it wrong i didnt square the bot terms

OpenStudy (anonymous):

okay wait

OpenStudy (anonymous):

25+16 =41 root 41 16+9 = 25 root 25

OpenStudy (anonymous):

so cos A = 32/root 41 * root 25 thats cos A = 0.999 A= arc cos 0.999 thats 2.5 .-. which is not an option

OpenStudy (anonymous):

Actually, I got 1.789910768 on my calculator which rounds off to 1.8 so you did fantastic aajugdar~!

OpenStudy (anonymous):

Thank you so much for helping me!

OpenStudy (anonymous):

yeahhh when i added 0.999512 gave me the same ^^

OpenStudy (anonymous):

Thank you for helping me, I hope that you have a great day. ^^

OpenStudy (anonymous):

welcome :) its night here though

OpenStudy (anonymous):

Whoops. Well, have a good night then. xD